Status Report

NOAA GOES SATOPS Morning Report 23 Jan 2001

By SpaceRef Editor
January 23, 2001
Filed under ,

SATOPS MORNING REPORT – January 23, 2001

POES

NOAA-16 CPU-A Anomaly Status:  Frist evident on January 16, the
backup on-board processor (CPU-A) has a timing problem within the CPU. 
CPU-A is currently processing instructions at a 10Hz speed, instead of
the appropriate 2 Hz speed.  The spacecraft is otherwise operating
normally and has been configured to eliminate to possibility of switching
to the anomalous processor.  The investigation has now shifted to
the recovery phase, as engineers from NOAA, NASA and the spacecraft contractor
are now investigating ways to change the clock speed via software within
the CPU.

NOAA-16 APT Anomaly Status:  From yesterday, the VHF Transmitter
1 (VTX1) was not turned off.  The transmitter was left in this configuration
due to the fact that the received AGCs observed of -98 dBm at WCDAS were
considered to be at least marginal and though no other independent verification
of good imagery on APT was available at that time, it was useful for further
testing.  Investigation and analysis continues.

NOAA-14 SEM:  Total Electron Detector (TED) shut down between 0330Z
and 0500Z last night.  Engineering advised.

NOAA-16:  Transferred L1A from rev 01666/Fairbanks to CEMSCS for
landmark verification.

NOAA-14:  Frame sync testing using begins tomorrow afternoon when
the new command schedule is activated.  Frame synch (F/S) 3 has been
configured for the test and is unusable as a backup.

NOAA-14 rev 31263 / F at 2352Z on January 22:  G1A noisy from spacecraft
with 93.9% recovered.  No retransmittal required.

NOAA-14 rev 31265 / F at 0315Z on January 23:  L1A noisy from spacecraft
with 99.2% recovered.  L1A not recoverable due to spacecraft recorder
sequence constraints.  G2B noisy from spacecraft with 99.6% recovered. 
100% G2B recovered on rev 31268.

DMSP

No change in the status of the DMSP spacecraft.

DMSP Launch Delayed for Undetermined Period of Time:  The chance
for a final attempt for launch today, January 23, was scrubbed because
of a problem detected with the satellite’s Inertial Measurement Unit (IMU)
that houses the attitude control gyroscopes.  Earlier Monday there
was hope that a switch to redundant battery power in this unit was due
to some disturbance during the booster leak investigation on Sunday night,
but the switch was not coincident with that activity.  Earlier problems
with an IMU cable and subsequent replacement, along with this latest incident,
forced Lockheed Martin to recommend a no-go for Tuesday.  With the
Titan rocket turnaround requirements and the scheduled Milstar launch support
by the Titan team, the estimate for rescheduling DMSP S-20 is about 70
days.  Thanks to everyone involved over the past few months in all
the preparation efforts for this mission.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

At 1306Z on January 22:  GOES-8 image frame break due to loss of
signal synch.

GOES operations were otherwise nominal over the past 24 hours.

 

Questions or comments to [email protected]

SpaceRef staff editor.